exactly three times as much to ODI tO GDP but a CD player does not contribute to GDP . PROBLEMS AND APPLICATIONS (50 marks) [1] Compute how much each of the following items is worth in terms of todays dollars using 177 as the price index for today. (15 marks) a. In 1926, the CPI was 17.7 and the price of a movie ticket was $0.25. b. In 1932, the CPI was 13.1 and a cook earned $15.00 a week. c. In 1943, the CPI was 17.4 and a gallon of gas cost $0.19. 21 Using separate graphs, demonstrate what happens to the money supply, money demand, the value of money, and the price level if: (15 marks) the Fed increases the money supply. a.
Answer #1

Solution:

A. $2.50 in today's dollars.
Working: CPI Current year/ CPI base year * Price index
= 177/17.7 * $0.25 = $2.50

B. $202.67 in today's dollars.
Working: 177/13.1* $15.00 = $202.67

C. $1.93 in today's dollars.
Working: 177/17.4 * $0.19= $1.93